### v1

@@ Abstract
The inference process in Large Language Models (LLMs) is often limited due to the absence of parallelism in the auto-regressive decoding process, resulting in most operations being restricted by the memory bandwidth of accelerators. While methods such as speculative decoding have been suggested to address this issue, their implementation is impeded by the challenges associated with acquiring and maintaining a separate draft model. In this paper, we present Medusa, an efficient method that augments LLM inference by adding extra decoding heads to predict multiple subsequent tokens in parallel. Using a tree-based attention mechanism, Medusa constructs multiple candidate continuations and verifies them simultaneously in each decoding step. By leveraging parallel processing, Medusa introduces only minimal overhead in terms of single-step latency while substantially reducing the number of decoding steps required.
We present two levels of fine-tuning procedures for Medusa to meet the needs of different use cases:
  • Medusa-1: Medusa is directly fine-tuned on top of a frozen backbone LLM, enabling lossless inference acceleration.
  • Medusa-2: Medusa is fine-tuned together with the backbone LLM, enabling better prediction accuracy of Medusa heads and higher speedup but needing a special training recipe that preserves the backbone model’s capabilities.
Moreover, we propose several extensions that improve or expand the utility of Medusa, including a self-distillation to handle situations where no training data is available and a typical acceptance scheme to boost the acceptance rate while maintaining generation quality.
We evaluate Medusa on models of various sizes and training procedures. Our experiments demonstrate that Medusa-1 can achieve over 2.2$\times$ speedup without compromising generation quality, while Medusa-2 further improves the speedup to 2.3-3.6$\times$. The code for this implementation is available at https://github.com/FasterDecoding/Medusa.
@@ 1 Introduction
The recent advancements in Large Language Models (LLMs) have demonstrated that the quality of language generation significantly improves with an increase in model size, reaching billions of parameters (Brown et al., 2020; Chowdhery et al., 2022; Zhang et al., 2022; Hoffmann et al., 2022; OpenAI, 2023; Google, 2023; Touvron et al., 2023). However, this growth has led to an increase in inference latency, which poses a significant challenge in practical applications. From a system perspective, LLM inference is predominantly memory-bound (Shazeer, 2019; Kim et al., 2023), with the main latency bottleneck stemming from accelerators’ memory bandwidth rather than arithmetic computations. This bottleneck is inherent to the sequential nature of auto-regressive decoding, where each forward pass requires transferring the complete model parameters from High-Bandwidth Memory (HBM) to the accelerator’s cache. This process, which generates only a single token, underutilizes the arithmetic computation potential of modern accelerators, leading to inefficiency.
@@
To address this, one approach to speed up LLM inference involves increasing the arithmetic intensity (the ratio of total floating-point operations (FLOPs) to total data movement) of the decoding process and reducing the number of decoding steps. In line with this idea, speculative decoding has been proposed (Leviathan et al., 2022; Chen et al., 2023; Xia et al., 2023; Miao et al., 2023). This method uses a smaller draft model to generate a sequence of tokens at each step, which is then refined by the original, larger model for acceptable continuation. However, obtaining an appropriate draft model remains challenging, and things become even harder when integrating the draft model into a distributed system (Chen et al., 2023).
@@
Instead of using a separate draft model to sequentially generate candidate outputs, in this paper, we revisit and refine the concept of using multiple decoding heads on top of the backbone model to expedite inference (Stern et al., 2018). We find that when applied effectively, this technique can overcome the challenges of speculative decoding, allowing for seamless integration into existing LLM systems. Specifically, we introduce Medusa, a method that enhances LLM inference by integrating additional decoding heads capable of concurrently predicting multiple tokens. These heads are fine-tuned in a parameter-efficient manner and can be added to any existing model. With no requirement for a new model, Medusa offers easy and automatic integration into current LLM systems, including those in distributed environments, ensuring a user-friendly experience.
@@
We further enhance Medusa with two key insights. Firstly, the current approach of generating a single candidate continuation at each decoding step leads to a restricted acceptance length and inefficient use of computational resources. To address this, we propose generating multiple candidate continuations using the Medusa heads and verifying them concurrently through a simple adjustment to the attention mask. Secondly, we can use the rejection sampling scheme similar to that used in speculative decoding to generate responses with the same distribution as the original model, but it is usually unnecessary for many LLM applications. Alternatively, we also introduce a typical acceptance scheme that selects reasonable candidates from the Medusa head outputs. We use temperature as a threshold to manage deviation from the original model’s predictions, providing an efficient alternative to the rejection sampling method. This approach effectively addresses its limitations, such as decreased speed at higher temperatures.
@@
Our experiments primarily focus on scenarios with a batch size of one, which is representative of the use case where LLMs are locally hosted for personal use [^1] It’s important to note that while Medusa can be seamlessly used in a batched inference setting, it requires additional engineering efforts to integrate into a serving engine like vLLM (Kwon et al., 2023). We are working on this and also welcome community contributions to help us. We test Medusa on models of varying sizes and training settings, including Vicuna-7B, 13B (trained with a public dataset), Vicuna-33B (Chiang et al., 2023) (trained with a private dataset [^2] Upon contacting the authors, this version is experimental and used some different data than Vicuna 7B and 13B.), and Zephyr-7B (trained with both supervised fine-tuning and alignment). Medusa can achieve a speedup of 2.3 to 3.6 times across different prompt types without compromising on the quality of generation.
Figure 1: Overview of Medusa. Medusa introduces multiple heads on top of the last hidden states of the LLM, enabling the prediction of several subsequent tokens in parallel (Section 3.1.1). For training Medusa heads, the original model is either frozen (Medusa-1, Section 3.2.1) or trained together (Medusa-2, Section 3.2.2) with Medusa heads. During inference, each head generates multiple top predictions for its designated position. These predictions are assembled into candidates, which are subsequently processed in parallel using a tree-based attention mechanism (Section 3.1.2). The final step is to verify the candidates and accept a continuation. Besides the standard rejection sampling scheme, a typical acceptance scheme (Section 3.3.1) can also be used here to select reasonable continuations, and the longest accepted candidate prefix will be used for the next decoding phase. The efficiency of the decoding process is enhanced by accepting more tokens simultaneously, thus reducing the number of required decoding steps.
2 Related Work
2.1 LLM Inference Acceleration
The inefficiency of Large Language Model (LLM) inference is primarily attributed to the memory-bound nature of the auto-regressive decoding process. Several methods have been proposed to alleviate this issue, improving inference latency and throughput. Traditionally, batch inference has been employed as a straightforward method to enhance arithmetic intensity and escape memory-bound limitations. However, with LLMs, both model parameters and the Key-Value (KV) cache consume substantial accelerator memory, hindering the utilization of large batch sizes. Existing methods to tackle this problem can be conceptually divided into two main categories: (1) Reducing memory consumption, thereby minimizing memory transfer overhead and enabling larger batch sizes, and (2) Minimizing the number of decoding steps to decrease latency directly.
Reducing KV Cache.
Quantization.
Speculative Decoding.
As an approach orthogonal to the aforementioned methods, speculative decoding (Leviathan et al., 2022; Chen et al., 2023) aims to execute several decoding steps in parallel, thus reducing the total number of steps required. This parallelization is realized by employing a smaller draft model to conjecture several subsequent words, which the LLMs then collectively evaluate and accept as appropriate. While resonating with non-autoregressive generation literature (Xiao et al., 2023b), this method is specifically tailored for LLMs to address the aforementioned inefficiency. Unlike previous works, we propose leveraging the original model to make predictions rather than introducing an additional draft model. This approach is more straightforward and seamlessly integrates into existing systems without the complexities of managing two models. Independently, Miao et al. (2023); Spector and Re (2023) propose the use of tree-structured attention to generate multiple candidates in parallel, where Miao et al. (2023) suggest employing an ensemble of models to propose candidates, and Spector and Re (2023) advocate adding another hierarchy for the draft model. After the first release of Medusa, we have seen many new works improving speculative decoding from the perspective of distillation (Liu et al., 2023; Zhou et al., 2023), making draft model training-free (He et al., 2023; Fu et al., 2023).
2.2 Sampling Scheme
3 Medusa
@@ 3.1.1 Medusa Heads
In speculative decoding, subsequent tokens are predicted by an auxiliary draft model. This draft model must be small yet effective enough to generate continuations that the original model will accept. Fulfilling these requirements is a challenging task, and existing approaches (Spector and Re, 2023; Miao et al., 2023) often resort to separately pre-training a smaller model. This pre-training process demands substantial additional computational resources. For example, in (Miao et al., 2023), a reported 275 NVIDIA A100 GPU hours were used. Additionally, separate pre-training can potentially create a distribution shift between the draft model and the original model, leading to continuations that the original model may not favor. Chen et al. (2023) have also highlighted the complexities of serving multiple models in a distributed environment.
To streamline and democratize the acceleration of LLM inference, we take inspiration from Stern et al. (2018) and introduce Medusa heads. These are additional decoding heads appended to the last hidden states of the original model. Specifically, given the original model’s last hidden states $h_{t}$ at position $t$, we add $K$ decoding heads to $h_{t}$. The $k$-th head is used to predict the token in the $(t+k+1)$-th position of the next tokens (the original language model head is used to predict the $(t+1)$-th position). The prediction of the $k$-th head is denoted as $p_{t}^{(k)}$, representing a distribution over the vocabulary, while the prediction of the original model is denoted as $p_{t}^{(0)}$. Following the approach of Stern et al. (2018), we utilize a single layer of feed-forward network with a residual connection for each head. We find that this simple design is sufficient to achieve satisfactory performance. The definition of the $k$-th head is outlined as:
$p_{t}^{(k)}=\text{softmax}\left(W_{2}^{(k)}\cdot\left(\text{SiLU}(W_{1}^{(k)}\cdot h_{t})+h_{t}\right)\right),\ \text{where }W_{2}^{(k)}\in\mathbb{R}^{d\times V},W_{1}^{(k)}\in\mathbb{R}^{d\times d}.$
We initialize $W_{1}^{(k)}$ identically to the original language model head, and $W_{2}^{(k)}$ to zero. This aligns the initial prediction of Medusa heads with that of the original model. The SiLU activation function (Elfwing et al., 2017) is employed following the Llama models (Touvron et al., 2023).
@@ 3.1.2 Tree Attention
Figure 2: Tree Attention Illustrated. This visualization demonstrates the use of tree attention to process multiple candidates concurrently. As exemplified, the top-2 predictions from the first Medusa head and the top-3 from the second result in a total of $2\times 3=6$ candidates. Each of these candidates corresponds to a distinct branch within the tree structure. To guarantee that each token only accesses its predecessors, we devise an attention mask that exclusively permits attention flow from the current token back to its antecedent tokens. The positional indices for positional encoding are adjusted in line with this structure.
This attention mechanism diverges from the traditional causal attention paradigm. Within this framework, only tokens from the same continuation are regarded as historical data. Drawing inspiration from the concept of embedding graph structures into attention as proposed in the graph neural network domain (Ying et al., 2021), we incorporate the tree structure into our attention mask, visualized in Figure 2. For a given $k$-th head, its top-$s_{k}$ predictions serve as the basis for candidate formation, where $s_{k}$ is a designated hyperparameter. These candidates are established by determining the Cartesian product of the top-$s_{k}$ predictions from each head. For instance, in Figure 2, with $s_{1}=2$ and $s_{2}=3$, each first head prediction can be succeeded by any prediction from the second head. This leads to a tree structure where $s_{k}$ branches exist at the $k$-th level (considering a virtual root as the 00-level, in practice, this 00-level is for the prediction of the language model head of the original model, which can be sampled independently). Within this tree, only a token’s predecessors are seen as historical context, and our attention mask ensures that the attention is only applied on a token’s predecessors. By employing this mask and properly setting the positional indices for positional encoding, we can process numerous candidates simultaneously without the need to expand the batch size. The cumulative number of new tokens is calculated as $\sum_{k=1}^{K}\prod_{i=1}^{k}s_{i}$.
@@ 3.2 Training Strategies
At the most basic level, we can train Medusa heads by freezing the backbone model and focusing solely on the Medusa heads. This approach is straightforward and requires minimal computational resources. However, training the backbone in conjunction with the Medusa heads can significantly enhance the accuracy of the Medusa heads. Depending on the computational resources and the specific requirements of the use case, we propose two levels of training strategies for Medusa heads.
In this section, we assume the availability of a training dataset that aligns with the target model’s output distribution. This could be the dataset used for Supervised Fine-Tuning (SFT) of the target model. We will discuss how to eliminate the need for such a dataset using a self-distillation approach in Section 3.3.2.
@@ 3.2.2 Medusa-2: Joint Training
  • Heads warmup: Noticing that at the beginning of training, the Medusa heads have a large loss, which leads to a large gradient and may distort the backbone model’s parameters. Following the idea from Kumar et al. (2022), we can employ a two-stage training process. In the first stage, we only train the backbone model as Medusa-1. In the second stage, we train the backbone model and Medusa heads together with a warmup strategy. Specifically, we first train the backbone model for a few epochs, then train the Medusa heads together with the backbone model. Besides this simple strategy, we can also use a more sophisticated warmup strategy by gradually increasing the weight $\lambda_{0}$ of the backbone model’s loss. We find both strategies work well in practice.
@@ 3.3.1 Typical Acceptance
In speculative decoding papers (Leviathan et al., 2022; Chen et al., 2023), authors employ rejection sampling to yield diverse outputs that align with the distribution of the original model. However, subsequent implementations (Joao Gante, 2023; Spector and Re, 2023) reveal that this sampling strategy results in diminished efficiency as the sampling temperature increases. Intuitively, this can be comprehended in the extreme instance where the draft model is the same as the original one. Here, when using greedy decoding, all output of the draft model will be accepted, therefore maximizing the efficiency. Conversely, rejection sampling introduces extra overhead, as the draft model and the original model are sampled independently. Even if their distributions align perfectly, the output of the draft model may still be rejected.
@@
However, in real-world scenarios, sampling from language models is often employed to generate diverse responses, and the temperature parameter is used merely to modulate the “creativity” of the response. Therefore, higher temperatures should result in more opportunities for the original model to accept the draft model’s output. We ascertain that it is typically unnecessary to match the distribution of the original model. Thus, we propose employing a typical acceptance scheme to select plausible candidates rather than using rejection sampling. This approach draws inspiration from truncation sampling studies (Hewitt et al., 2022) (refer to Section 2 for an in-depth explanation). Our objective is to choose candidates that are typical, meaning they are not exceedingly improbable to be produced by the original model. We use the prediction probability from the original model as a natural gauge for this and establish a threshold based on the prediction distribution to determine acceptance. Specifically, given $x_{1},x_{2},\cdots,x_{n}$ as context, when evaluating the candidate sequence $(x_{n+1},x_{n+2},\cdots,x_{n+T+1})$ (composed by top predictions of the original language model head and Medusa heads), we consider the condition
@@
$p_{\text{original}}(x_{n+k}|x_{1},x_{2},\cdots,x_{n+k-1})>\min\left(\epsilon,\delta\exp\left(-H(p_{\text{original}}(\cdot|x_{1},x_{2},\cdots,x_{n+k-1}))\right)\right),$
@@
where $H(\cdot)$ denotes the entropy function, and $\epsilon,\delta$ are hyperparameters. This criterion is adapted from Hewitt et al. (2022) and rests on two observations: (1) tokens with relatively high probability are meaningful, and (2) when the distribution’s entropy is high, various continuations may be deemed reasonable. During decoding, every candidate is evaluated using this criterion, and a prefix of the candidate is accepted if it satisfies the condition. To guarantee the generation of at least one token at each step, we apply greedy decoding for the first token and unconditionally accept it while employing typical acceptance for subsequent tokens. The final prediction for the current step is determined by the longest accepted prefix among all candidates.

### v2

@@ Abstract
Large Language Models (LLMs) employ auto-regressive decoding that requires sequential computation, with each step reliant on the previous one’s output. This creates a bottleneck as each step necessitates moving the full model parameters from High-Bandwidth Memory (HBM) to the accelerator’s cache. While methods such as speculative decoding have been suggested to address this issue, their implementation is impeded by the challenges associated with acquiring and maintaining a separate draft model. In this paper, we present Medusa, an efficient method that augments LLM inference by adding extra decoding heads to predict multiple subsequent tokens in parallel. Using a tree-based attention mechanism, Medusa constructs multiple candidate continuations and verifies them simultaneously in each decoding step. By leveraging parallel processing, Medusa substantially reduces the number of decoding steps required. We present two levels of fine-tuning procedures for Medusa to meet the needs of different use cases: Medusa-1: Medusa is directly fine-tuned on top of a frozen backbone LLM, enabling lossless inference acceleration. Medusa-2: Medusa is fine-tuned together with the backbone LLM, enabling better prediction accuracy of Medusa heads and higher speedup but needing a special training recipe that preserves the model’s capabilities. Moreover, we propose several extensions that improve or expand the utility of Medusa, including a self-distillation to handle situations where no training data is available and a typical acceptance scheme to boost the acceptance rate while maintaining generation quality. We evaluate Medusa on models of various sizes and training procedures. Our experiments demonstrate that Medusa-1 can achieve over 2.2$\times$ speedup without compromising generation quality, while Medusa-2 further improves the speedup to 2.3-2.8$\times$.
Machine Learning, ICML
@@ 1 Introduction
The recent advancements in Large Language Models (LLMs) have demonstrated that the quality of language generation significantly improves with an increase in model size, reaching billions of parameters (Brown et al., 2020; Chowdhery et al., 2022; Zhang et al., 2022; Hoffmann et al., 2022; OpenAI, 2023; Google, 2023; Touvron et al., 2023). However, this growth has led to an increase in inference latency, which poses a significant challenge in practical applications. From a system perspective, LLM inference is predominantly memory-bandwidth-bound (Shazeer, 2019; Kim et al., 2023), with the main latency bottleneck stemming from accelerators’ memory bandwidth rather than arithmetic computations. This bottleneck is inherent to the sequential nature of auto-regressive decoding, where each forward pass requires transferring the complete model parameters from High-Bandwidth Memory (HBM) to the accelerator’s cache. This process, which generates only a single token, underutilizes the arithmetic computation potential of modern accelerators, leading to inefficiency.
@@
To address this, one approach to speed up LLM inference involves increasing the arithmetic intensity (the ratio of total floating-point operations (FLOPs) to total data movement) of the decoding process and reducing the number of decoding steps. In line with this idea, speculative decoding has been proposed (Leviathan et al., 2022; Chen et al., 2023; Xia et al., 2023; Miao et al., 2023). This method uses a smaller draft model to generate a token sequence, which is then refined by the original, larger model for acceptable continuation. However, obtaining an appropriate draft model remains challenging, and it’s even harder to integrate the draft model into a distributed system (Chen et al., 2023).
@@
Instead of using a separate draft model to sequentially generate candidate outputs, in this paper, we revisit and refine the concept of using multiple decoding heads on top of the backbone model to expedite inference (Stern et al., 2018). We find that when applied effectively, this technique can overcome the challenges of speculative decoding, allowing for seamless integration into existing LLM systems. Specifically, we introduce Medusa, a method that enhances LLM inference by integrating additional decoding heads to concurrently predict multiple tokens. These heads are fine-tuned in a parameter-efficient manner and can be added to any existing model. With no requirement for a draft model, Medusa offers easy integration into current LLM systems, including those in distributed environments, ensuring a user-friendly experience.
@@
We further enhance Medusa with two key insights. Firstly, the current approach of generating a single candidate continuation at each decoding step leads to inefficient use of computational resources. To address this, we propose generating multiple candidate continuations using the Medusa heads and verifying them concurrently through a simple adjustment to the attention mask. Secondly, we can reuse the rejection sampling scheme as used in speculative decoding (Leviathan et al., 2022; Chen et al., 2023) to generate consistent responses with the same distribution as the original model. However, it cannot further enhance the acceleration rate. Alternatively, we introduce a typical acceptance scheme that selects reasonable candidates from the Medusa head outputs. We use temperature as a threshold to manage deviation from the original model’s predictions, providing an efficient alternative to the rejection sampling method. Our results suggest that the proposed typical acceptance scheme can accelerate the decoding speed further while maintaining a similar generation quality.
@@
Our experiments primarily focus on scenarios with a batch size of one, which is representative of the use case where LLMs are locally hosted for personal use. We test Medusa on models of varying sizes and training settings, including Vicuna-7B, 13B (trained with a public dataset), Vicuna-33B (Chiang et al., 2023) (trained with a private dataset [^1] Upon contacting the authors, this version is experimental and used some different data than Vicuna 7B and 13B.), and Zephyr-7B (trained with both supervised fine-tuning and alignment). Medusa can achieve a speedup of 2.3 to 2.8 times across different prompt types without compromising on the quality of generation.
Figure 1: Medusa introduces multiple heads on top of the last hidden states of the LLM, enabling the prediction of several subsequent tokens in parallel (Section 2.1.1). During inference, each head generates multiple top predictions for its designated position. These predictions are assembled into candidates, which are processed in parallel using a tree-based attention mechanism (Section 2.1.2). The final step is to verify the candidates and accept a continuation. Besides the standard rejection sampling scheme, a typical acceptance scheme (Section 2.3.1) can also be used here to select reasonable continuations, and the longest accepted candidate prefix will be used for the next decoding phase.
2 Methodology
@@ 2.1.1 Medusa Heads
In speculative decoding, subsequent tokens are predicted by an auxiliary draft model. This draft model must be small yet effective enough to generate continuations that the original model will accept. Fulfilling these requirements is a challenging task, and existing approaches (Spector & Re, 2023; Miao et al., 2023) often resort to separately pre-training a smaller model. This pre-training process demands substantial additional computational resources. For example, in (Miao et al., 2023), a reported 275 NVIDIA A100 GPU hours were used. Additionally, separate pre-training can potentially create a distribution shift between the draft model and the original model, leading to continuations that the original model may not favor. Chen et al. (2023) have also highlighted the complexities of serving multiple models in a distributed environment.
To streamline and democratize the acceleration of LLM inference, we take inspiration from Stern et al. (2018), which utilizes parallel decoding for tasks such as machine translation and image super-resolution. Medusa heads are additional decoding heads appended to the last hidden states of the original model. Specifically, given the original model’s last hidden states $h_{t}$ at position $t$, we add $K$ decoding heads to $h_{t}$. The $k$-th head is used to predict the token in the $(t+k+1)$-th position of the next tokens (the original language model head is used to predict the $(t+1)$-th position). The prediction of the $k$-th head is denoted as $p_{t}^{(k)}$, representing a distribution over the vocabulary, while the prediction of the original model is denoted as $p_{t}^{(0)}$. Following the approach of Stern et al. (2018), we utilize a single layer of feed-forward network with a residual connection for each head. We find that this simple design is sufficient to achieve satisfactory performance. The definition of the $k$-th head is outlined as:
$p_{t}^{(k)}=\text{softmax}\left(W_{2}^{(k)}\cdot\left(\text{SiLU}(W_{1}^{(k)}\cdot h_{t})+h_{t}\right)\right),$
$\text{where }W_{2}^{(k)}\in\mathbb{R}^{d\times V},W_{1}^{(k)}\in\mathbb{R}^{d\times d}.$
$d$ is the output dimension of the LLM’s last hidden layer and $V$ is the vocabulary size. We initialize $W_{2}^{(k)}$ identically to the original language model head, and $W_{1}^{(k)}$ to zero. This aligns the initial prediction of Medusa heads with that of the original model. The SiLU activation function (Elfwing et al., 2017) is employed following the Llama models (Touvron et al., 2023).
@@ 2.1.2 Tree Attention
Figure 2: We demonstrates the use of tree attention to process multiple candidates concurrently. As exemplified, the top-2 predictions from the first Medusa head and the top-3 from the second result in a total of $2\times 3=6$ candidates. Each of these candidates corresponds to a distinct branch within the tree structure. To guarantee that each token only accesses its predecessors, we devise an attention mask that exclusively permits attention flow from the current token back to its antecedent tokens. The positional indices for positional encoding are adjusted in line with this structure.
This attention mechanism diverges from the traditional causal attention paradigm. Within this framework, only tokens from the same continuation are regarded as historical data. Drawing inspiration from the concept of embedding graph structures into attention as proposed in the graph neural network domain (Ying et al., 2021), we incorporate the tree structure into our attention mask, visualized in Figure 2. Remarkably, similar ideas have also been explored in independent works like Miao et al. (2023); Spector & Re (2023), where they follow a bottom-up approach and construct the tree by merging multiple candidates generated by a draft model. In our method, we instead take a top-down approach to build the tree thanks to the structure of candidates generated by Medusa heads. For a given $k$-th head, its top-$s_{k}$ predictions serve as the basis for candidate formation, where $s_{k}$ is a designated hyperparameter. These candidates are established by determining the Cartesian product of the top-$s_{k}$ predictions from each head. For instance, in Figure 2, with $s_{1}=2$ and $s_{2}=3$, each first head prediction can be succeeded by any prediction from the second head. This leads to a tree structure where $s_{k}$ branches exist at the $k$-th level (considering a virtual root as the 00-level, in practice, this 00-level is for the prediction of the language model head of the original model, which can be sampled independently). Within this tree, only a token’s predecessors are seen as historical context, and our attention mask ensures that the attention is only applied on a token’s predecessors. By employing this mask and properly setting the positional indices for positional encoding, we can process numerous candidates simultaneously without the need to expand the batch size. The cumulative number of new tokens is calculated as $\sum_{k=1}^{K}\prod_{i=1}^{k}s_{i}$.
@@ 2.2 Training Strategies
At the most basic level, we can train Medusa heads by freezing the backbone model and fine-tuning Medusa heads. However, training the backbone in conjunction with the Medusa heads can significantly enhance the accuracy of the Medusa heads. Depending on the computational resources and the specific reqirements of the use case, we propose two levels of training strategies for Medusa heads.
In this section, we assume the availability of a training dataset that aligns with the target model’s output distribution. This could be the dataset used for Supervised Fine-Tuning (SFT) of the target model. We will discuss eliminating the need for such a dataset using a self-distillation approach in Section 2.3.2.
@@ 2.2.2 Medusa-2: Joint Training
  • Heads warmup: Noticing that at the beginning of training, the Medusa heads have a large loss, which leads to a large gradient and may distort the backbone model’s parameters. Following the idea from Kumar et al. (2022), we can employ a two-stage training process. In the first stage, we only train the Medusa heads as Medusa-1. In the second stage, we train the backbone model and Medusa heads together with a warmup strategy. Specifically, we first train the backbone model for a few epochs, then train the Medusa heads together with the backbone model. Besides this simple strategy, we can also use a more sophisticated warmup strategy by gradually increasing the weight $\lambda_{0}$ of the backbone model’s loss. We find both strategies work well in practice.
@@
2.2.3 How to Select the Number of Heads
Empirically, we found that five heads are sufficient at most. Therefore, we recommend training with five heads and referring to the strategy described in Section 2.3.3 to determine the optimal configuration of the tree attention. With optimized tree attention, sometimes three or four heads may be enough for inference. In this case, we can ignore the redundant heads without overhead.
@@ 2.3.1 Typical Acceptance
In speculative decoding papers (Leviathan et al., 2022; Chen et al., 2023), authors employ rejection sampling to yield diverse outputs that align with the distribution of the original model. However, subsequent implementations (Joao Gante, 2023; Spector & Re, 2023) reveal that this sampling strategy results in diminished efficiency as the sampling temperature increases. Intuitively, this can be comprehended in the extreme instance where the draft model is the same as the original one: Using greedy decoding, all output of the draft model will be accepted, therefore maximizing the efficiency. Conversely, rejection sampling introduces extra overhead, as the draft model and the original model are sampled independently. Even if their distributions align perfectly, the output of the draft model may still be rejected.
@@
However, in real-world scenarios, sampling from language models is often employed to generate diverse responses, and the temperature parameter is used merely to modulate the “creativity” of the response. Therefore, higher temperatures should result in more opportunities for the original model to accept the draft model’s output. We ascertain that it is typically unnecessary to match the distribution of the original model. Thus, we propose employing a typical acceptance scheme to select plausible candidates rather than using rejection sampling. This approach draws inspiration from truncation sampling studies (Hewitt et al., 2022) (refer to Appendix A for an in-depth explanation). Our objective is to choose candidates that are typical, meaning they are not exceedingly improbable to be produced by the original model. We use the prediction probability from the original model as a natural gauge for this and establish a threshold based on the prediction distribution to determine acceptance. Specifically, given $x_{1},x_{2},\cdots,x_{n}$ as context, when evaluating the candidate sequence $(x_{n+1},x_{n+2},\cdots,x_{n+K+1})$ (composed by top predictions of the original language model head and Medusa heads), we consider the condition
$p_{\text{original}}(x_{n+k}|x_{1},x_{2},\cdots,x_{n+k-1})>$
@@
$\min\left(\epsilon,\delta\exp\left(-H(p_{\text{original}}(\cdot|x_{1},x_{2},\cdots,x_{n+k-1}))\right)\right),$
@@
where $H(\cdot)$ denotes the entropy function, and $\epsilon,\delta$ are the hard threshold and the entropy-dependent threshold respectively. This criterion is adapted from Hewitt et al. (2022) and rests on two observations: (1) tokens with relatively high probability are meaningful, and (2) when the distribution’s entropy is high, various continuations may be deemed reasonable. During decoding, every candidate is evaluated using this criterion, and a prefix of the candidate is accepted if it satisfies the condition. To guarantee the generation of at least one token at each step, we apply greedy decoding for the first token and unconditionally accept it while employing typical acceptance for subsequent tokens. The final prediction for the current step is determined by the longest accepted prefix among all candidates.
@@
Empirically, we verify that typical acceptance can achieve a better speedup while maintaining a similar generation quality as shown in Figure 5.